OpenStudy (anonymous):

Using complete sentences, explain the difference between the equations of a circle and an ellipse.

OpenStudy (turingtest):

\[x^2+y^2=r^2\to\frac{x^2}{r^2}+\frac{y^2}{r^2}=1\]is a circle\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]is an ellipse r, a, and b are constants. What differences are there between the two equations?